NOAA Office of Satellite and Product Operations

NOAA 19 HIRS Subsystem Summary

Component Description Status
CHANNEL 01 Spectral Channel #01, 14.95 micro meters Orange
CHANNEL 02 Spectral Channel #02, 14.71 micro meters Orange
CHANNEL 03 Spectral Channel #03, 14.49 micro meters Orange
CHANNEL 04 Spectral Channel #04, 14.22 micro meters Orange
CHANNEL 05 Spectral Channel #05, 13.97 micro meters Orange
CHANNEL 06 Spectral Channel #06, 13.64 micro meters Orange
CHANNEL 07 Spectral Channel #07, 13.35 micro meters Orange
CHANNEL 08 Spectral Channel #08, 11.11 micro meters Orange
CHANNEL 09 Spectral Channel #09, 9.71 micro meters Orange
CHANNEL 10 Spectral Channel #10, 12.47 micro meters Orange
CHANNEL 11 Spectral Channel #11, 7.33 micro meters Orange
CHANNEL 12 Spectral Channel #12, 6.52 micro meters Orange
CHANNEL 13 Spectral Channel #13, 4.57 micro meters Green
CHANNEL 14 Spectral Channel #14, 4.52 micro meters Orange
CHANNEL 15 Spectral Channel #15, 4.47 micro meters Green
CHANNEL 16 Spectral Channel #16, 4.45 micro meters Green
CHANNEL 17 Spectral Channel #17, 4.13 micro meters Green
CHANNEL 18 Spectral Channel #18, 4.00 micro meters Green
CHANNEL 19 Spectral Channel #19, 3.76 micro meters Green
CHANNEL 20 Spectral Channel #20, 0.69 micro meters Green
ELECTRONICS Instrument Electronics Green
FILTER WHEEL Instrument Filter Wheel Motors Orange
SCAN MOTOR Instrument Scan Motor Green

Subsystem History:

Date Time Component Description
03/16/2015   CHANNEL 14 Channel 14 radiometric and warm target counts went from degraded to saturated from 10:55UTC until 12:10UTC
01/26/201517:00FILTER WHEELFilter wheel returned to nominal values
01/26/201511:15FILTER WHEELFilter wheel became unsynced as the filter wheel current spiked multiple times reaching a value of about 377.2mA
01/26/20151:00FILTER WHEELFilter wheel returned to nominal values
01/25/20156:00FILTER WHEELFilter wheel returned to nominal values
01/25/20154:30FILTER WHEELFilter wheel became unsynced as the filter wheel current spiked reaching a value of about 390mA
01/25/2015FILTER WHEELFilter wheel became unsynced as the filter wheel current spiked reaching a value of about 387.7mA
01/23/201522:00FILTER WHEELFilter wheel returned to nominal values
01/23/201518:30FILTER WHEELFilter wheel became unsynced as the filter wheel current spiked multiple times reaching a value of 386mA
01/19/20155:00FILTER WHEELFilter wheel returned to nominal values
01/19/20152:30FILTER WHEELFilter wheel became unsynced as the filter wheel current spiked multiple times reaching a value of 379.2mA.
01/18/20159:30CHANNEL 14Channel 14 radiometric counts returned to nominal values
01/18/20152:10CHANNEL 14Channel 14 radiometric counts degrade and became saturated, briefly returned from 5:10 until 6:10, then saturated again.
01/18/201519:00CHANNEL 14Channel 14 radiometric counts returned to nominal values
01/18/201519:00FILTER WHEELFilter wheel returned to nominal values
01/18/201517:30CHANNEL 14Channel 14 radiometric counts degrade and became saturated
01/18/20151:15FILTER WHEELFilter wheel became unsynced as the filter wheel current spiked numerous times between 1:15 and 8:30 reaching the maximum value for the day of 401mA. Later the filter wheel current spiked several more times starting at 15:45.
01/13/201522:00CHANNEL 14Channel 14 radiometric counts returned to nominal values
01/13/201522:00FILTER WHEELFilter wheel returned to nominal operations
01/13/201520:00CHANNEL 14Channel 14 radiometric counts degrade and became saturated
01/13/201519:00FILTER WHEELFilter wheel became unsynced as the filter wheel current spiked reaching a value of 379.2mA.
12/23/20146:00CHANNEL 14Channel 14 radiometric counts degrade and became saturated
12/23/20144:50FILTER WHEELFilter wheel became unsynced as the filter wheel current spiked reaching a value of 374.8mA., Returned briefly at 8:00 UTC, before spiking again at 12:15
12/23/201414:00FILTER WHEELFilter wheel returned to nominal operations
12/23/201410:45CHANNEL 14Channel 14 radiometric counts returned to nominal values
12/22/20145:15FILTER WHEELFilter wheel became unsynced as the filter wheel current spiked reaching a value of 375.3mA.
12/22/201410:45FILTER WHEELFilter wheel returned to nominal operations
12/17/20149:00CHANNEL 14Channel 14 radiometric counts degrade and became saturated
12/17/20148:20FILTER WHEELFilter wheel became unsynced as the filter wheel current spiked reaching a value of 372.8mA.
12/17/201411:00CHANNEL 14Channel 14 radiometric counts returned to nominal values
12/17/201411:00FILTER WHEELFilter wheel returned to nominal operations
12/13/201419:15CHANNEL 14Channel 14 radiometric counts unsaturated
12/13/201419:15FILTER WHEELFilter wheel returned to nominal operations
12/13/201417:15CHANNEL 14Channel 14 radiometric counts saturated
12/13/201416:30FILTER WHEELFilter wheel became unsynced as the filter wheel current spiked reaching a value of 384mA.
12/09/201400:30FILTER WHEELFilter wheel returned to nominal operations
12/08/201422:30FILTER WHEELFilter wheel became unsynced as the filter wheel current spiked reaching a value of 387mA.
12/03/201416:30FILTER WHEELFilter wheel returned to nominal operations
12/03/201414:00FILTER WHEELFilter wheel became unsynced as the filter wheel current spiked reaching a value of 394.8mA.
12/01/20147:00CHANNEL 14Channel 14 radiometric counts saturated
12/01/20144:30FILTER WHEELFilter wheel became unsynced as the filter wheel current spiked reaching a value of 406mA.
12/01/201410:30CHANNEL 14Channel 14 radiometric counts unsaturated
12/01/201410:30FILTER WHEELFilter wheel returned to nominal operations
11/29/20145:00CHANNEL 14Channel 14 radiometric counts unsaturated
11/29/20145:00FILTER WHEELFilter wheel returned to nominal operations
11/29/20142:45CHANNEL 14Channel 14 radiometric counts saturated
11/29/20142:00FILTER WHEELFilter wheel became unsynced as the filter wheel current spiked reaching a value of 400mA.
11/24/20149:30CHANNEL 14Channel 14 radiometric counts unsaturated
11/24/20149:30FILTER WHEELFilter wheel returned to nominal operations
11/24/20147:30CHANNEL 14Channel 14 radiometric counts saturated
11/24/20146:45FILTER WHEELFilter wheel became unsynced as the filter wheel current spiked reaching a value of 396.8mA.
11/19/201417:00CHANNEL 14Channel 14 radiometric counts unsaturated
11/19/201416:30FILTER WHEELFilter wheel returned to nominal operations
11/19/201413:30CHANNEL 14Channel 14 radiometric counts saturated
11/19/201412:15FILTER WHEELFilter wheel became unsynced as the filter wheel current spiked reaching a value of 387mA.
11/16/201423:00CHANNEL 14Channel 14 radiometric counts unsaturated
11/16/201422:00FILTER WHEELFilter wheel returned to nominal operations
11/16/201420:30CHANNEL 14Channel 14 radiometric counts saturated
11/16/201420:30FILTER WHEELFilter wheel became unsynced as the filter wheel current spiked reaching a value of 384mA.
11/09/20146:00FILTER WHEELFilter wheel returned to nominal operations
11/09/20145:25FILTER WHEELFilter wheel became unsynced as the filter wheel current spiked reaching a value of 392mA. Channel 14 only exhibited a few bad points during the increased current.
10/23/20148:00FILTER WHEELFilter wheel returned to nominal operations
10/22/201421:30CHANNEL 14Channel 14 radiometric counts unsaturated
10/22/201412:25CHANNEL 14Channel 14 radiometric counts saturated
10/22/201411:15FILTER WHEELFilter wheel became unsynced as the filter wheel current spiked reaching a value of 384mA. It lasted through 23 Oct 2014 (14/296) 1:30 UTC when the motor returned to nominal levels for a few hours then spiked again at 6:10 UTC.
10/08/201421:30CHANNEL 14Channel 14 radiometric counts unsaturated
10/08/201421:30FILTER WHEELFilter wheel returned to nominal operations
10/08/201415:30CHANNEL 14Channel 14 radiometric counts saturated
10/08/201415:15FILTER WHEELFilter wheel became unsynced as the filter wheel current spiked reaching a value of 385.2mA
10/04/20145:35CHANNEL 14Channel 14 radiometric counts unsaturated
10/04/20145:30FILTER WHEELFilter wheel returned to nominal operations
10/04/201400:30CHANNEL 14Channel 14 radiometric counts saturated
10/03/201423:55FILTER WHEELFilter wheel became unsynced as the filter wheel current spiked reaching a value of 375.4mA
09/25/201414:30FILTER WHEELFilter wheel returned to nominal operations
09/25/201412:45FILTER WHEELFilter wheel became unsynced as the filter wheel current spiked reaching a value of 375.4mA
09/20/20146:45FILTER WHEELFilter wheel returned to nominal operations
09/20/20145:20FILTER WHEELFilter wheel became unsynced as the filter wheel current spiked reaching a value of 398.7mA
09/18/201423:00FILTER WHEELFilter wheel returned to nominal operations
09/18/201420:30FILTER WHEELFilter wheel became unsynced as the filter wheel current spiked reaching a value of 379.7mA
09/14/201421:00FILTER WHEELFilter wheel returned to nominal operations
09/14/201416:45FILTER WHEELFilter wheel became unsynced as the filter wheel current spiked reaching a value of 391mA
09/13/201423:00FILTER WHEELFilter wheel returned to nominal operations
09/13/201420:45FILTER WHEELFilter wheel became unsynced as the filter wheel current spiked reaching a value of 391mA.
09/10/20145:00CHANNEL 14Channel 14 radiometric counts saturated, returned briefly from 9UTC to 13:00UTC before saturating again for roughly another 7 hours.
09/10/20143:00FILTER WHEELFilter wheel became unsynced as the filter wheel current spiked reaching a value of 375.2mA, The current renturned to a nominal level for about 5 hours (7:30-11:30UTC) before another current surge occured. The filter sync continued to waiver for the next 12 hours.
09/10/201423:30FILTER WHEELFilter wheel returned to nominal operations
09/10/201420:00CHANNEL 14Channel 14 radiometric counts unsaturated
09/06/201413:00FILTER WHEELFilter wheel returned to nominal operations
09/06/201410:15FILTER WHEELFilter wheel became unsynced as the filter wheel current spiked briefly reaching a value of 375.2mA
08/28/20142:20CHANNEL 14Channel 14 radiometric counts unsaturated
08/28/20142:00FILTER WHEELFilter wheel returned to nominal operations
08/27/201421:10CHANNEL 14Channel 14 radiometric counts saturated
08/27/201420:45 FILTER WHEELFilter wheel became unsynced as the filter wheel current spiked briefly reaching a value of 379.2mA
08/08/201419:05CHANNEL 14Channel 14 radiometric counts unsaturated
08/08/201419:05FILTER WHEELFilter wheel returned to nominal operations
08/08/201413:00CHANNEL 14Channel 14 radiometric counts saturated
08/08/201412:20FILTER WHEELFilter wheel became unsynced as the filter wheel current spiked briefly reaching a value of 438mA
07/25/20145:20CHANNEL 14Channel 14 radiometric counts saturated
07/25/20144:37FILTER WHEELFilter wheel became unsynced as the filter wheel current spiked briefly reaching a value of 379.2mA
07/25/201411:45CHANNEL 14Channel 14 radiometric counts unsaturated
07/25/201411:45FILTER WHEELFilter wheel returned to nominal operations
06/28/20145:00FILTER WHEELFilter wheel returned to nominal operations
06/28/20144:35CHANNEL 14Channel 14 radiometric counts unsaturated
06/27/201418:15CHANNEL 14Channel 14 radiometric counts saturated
06/27/201417:28FILTER WHEELFilter wheel became unsynced as the filter wheel current spiked briefly reaching a value of 405mA
06/20/20149:15CHANNEL 14Channel 14 radiometric counts unsaturated
06/20/20149:00FILTER WHEELFilter wheel returned to nominal operations
06/20/20145:00CHANNEL 14Channel 14 radiometric counts saturated
06/20/20144:20FILTER WHEELFilter wheel became unsynced as the filter wheel current spiked briefly reaching a value of 380mA
06/11/201405:55CHANNEL 14Channel 14 radiometric counts became unsaturated
06/11/201404:30FILTER WHEELFilter wheel returned to nominal operations
06/11/201401:07CHANNEL 14Channel 14 radiometric counts became saturated
06/11/201400:20FILTER WHEELFilter wheel became unsynced as the filter wheel current spiked briefly reaching a value of 398mA
06/01/201420:15CHANNEL 14Channel 14 radiometric counts became saturated
06/01/201419:25FILTER WHEELFilter wheel became unsynced as the filter wheel current spiked briefly reaching a value of 381mA
06/01/201400:55CHANNEL 14Channel 14 radiometric counts became unsaturated
06/01/201400:15FILTER WHEELFilter wheel returned to nominal operations
05/18/20142:15FILTER WHEELFilter wheel became unsynced as the filter wheel current spiked briefly reaching a value of 373mA
05/18/201407:05CHANNEL 14Channel 14 radiometric counts became unsaturated
05/18/201406:45FILTER WHEELFilter wheel returned to nominal operations
05/18/201402:45CHANNEL 14Channel 14 radiometric counts became saturated
05/18/201402:05CHANNEL 14Channel 14 radiometric counts became unsaturated
05/18/201401:15FILTER WHEELFilter wheel returned to nominal operations
05/17/201421:47CHANNEL 14Channel 14 radiometric counts became saturated
05/17/201420:50FILTER WHEELFilter wheel became unsynced as the filter wheel current spiked briefly reaching a value of 382mA
01/28/2014 03:40:00 FILTER WHEEL HIRS Filter Wheel began experiencing current surges 1/28/2014. Presently, the FW is experiencing these surges every 5-10 days lasting for a few hours causing all Channel radiometric patterns to shift and Channel 14's radiometrics to saturate.
11/30/2013 07:00:00 CHANNEL 14 Due to drop in Filter Wheel motor current, Channel 14 radiometrics counts unsaturated.
11/20/2013 13:30:00 CHANNEL 14 Channel 14 radiometrics return to saturation levels.
11/17/2013 13:30:00 CHANNEL 14 Due to drop in Filter Wheel motor current, Channel 14 radiometrics became unsaturated.
10/9/2013 09:00:00 CHANNEL 14 Channel became unsaturated; Internal Target counts began saturating around JDAY 284, by JDAY 286 most radioimetric counts are saturated again. FW motor current dropped which allowed Chanel 14 radiometrics to unsaturate for a few days.
09/5/2013 07:00:00 CHANNEL 14 Channel 14 Warm Target, Space View and Earth View radiometrics continue to have periods of nominal radiometric counts but are mostly staying at saturated levels.
09/5/2013 00:00:00 CHANNEL 14 Channel 14 Warm Target, Space View and Earth View radiometrics continue to have periods of nominal radiometric counts but are mostly staying at saturated levels.
07/26/2013 00:00:00 CHANNEL 14 Channel 14 Warm Target, Space View and Earth View radiometrics saturated for a few days before
07/12/2013 00:00:00   Due to User's inability to use the HIRS data the instrurment is declared Orange
07/11/2013 00:00:00 CHANNEL 06 Increased IR Channel Noise has been reduced
07/11/2013 00:00:00 CHANNEL 01 Increased IR Channel Noise has been reduced
07/11/2013 00:00:00 CHANNEL 02 Increased IR Channel Noise has been reduced
07/11/2013 00:00:00 CHANNEL 03 Increased IR Channel Noise has been reduced
07/11/2013 00:00:00 CHANNEL 05 Increased IR Channel Noise has been reduced
07/11/2013 00:00:00 CHANNEL 07 Increased IR Channel Noise has been reduced
07/11/2013 00:00:00 CHANNEL 08 Increased IR Channel Noise has been reduced
07/11/2013 00:00:00 CHANNEL 09 Increased IR Channel Noise has been reduced
07/11/2013 00:00:00 CHANNEL 10 Increased IR Channel Noise has been reduced
07/11/2013 00:00:00 CHANNEL 11 Increased IR Channel Noise has been reduced
07/11/2013 00:00:00 CHANNEL 12 Increased IR Channel Noise has been reduced
07/11/2013 00:00:00   Increased IR Channel 1-12 Noise has been reduced
07/11/2013 00:00:00 FILTER WHEEL Increased erratic current has improved
07/11/2013 00:00:00 CHANNEL 04 Increased IR Channel Noise has been reduced
07/10/2013 18:20:02 FILTER WHEEL Filter housing heat ON
07/10/2013 18:19:28 FILTER WHEEL Filter Motor switched to High Power
07/02/2013 23:00:00 CHANNEL 01 Increased IR Channel Noise
07/02/2013 23:00:00 CHANNEL 09 Increased IR Channel Noise
07/02/2013 23:00:00 CHANNEL 05 Increased IR Channel Noise
07/02/2013 23:00:00 CHANNEL 02 Increased IR Channel Noise
07/02/2013 23:00:00 CHANNEL 03 Increased IR Channel Noise
07/02/2013 23:00:00 CHANNEL 08 Increased IR Channel Noise
07/02/2013 23:00:00 CHANNEL 11 Increased IR Channel Noise
07/02/2013 23:00:00 CHANNEL 12 Increased IR Channel Noise
07/02/2013 23:00:00 CHANNEL 06 Increased IR Channel Noise
07/02/2013 23:00:00 CHANNEL 07 Increased IR Channel Noise
07/02/2013 23:00:00 FILTER WHEEL Increased erratic behavior of the current was observed
07/02/2013 23:00:00 CHANNEL 04 Increased IR Channel Noise
07/02/2013 23:00:00 CHANNEL 10 Increased IR Channel Noise
09/08/2011 22:06:50 ELECTRONICS Scan Mirror Temperature sensor returned to nominal behavior.
09/02/2011 01:20:00 ELECTRONICS Scan Mirror Temperature sensor failed; no impact to products
02/22/2009 12:30:00 CHANNEL 03 HIRS Electronics ON
02/22/2009 12:30:00 CHANNEL 10 HIRS Electronics ON
02/22/2009 12:30:00 CHANNEL 09 HIRS Electronics ON
02/22/2009 12:30:00 CHANNEL 08 HIRS Electronics ON
02/22/2009 12:30:00 CHANNEL 07 HIRS Electronics ON
02/22/2009 12:30:00 CHANNEL 06 HIRS Electronics ON
02/22/2009 12:30:00 CHANNEL 11 HIRS Electronics ON
02/22/2009 12:30:00 CHANNEL 04 HIRS Electronics ON
02/22/2009 12:30:00 CHANNEL 17 HIRS Electronics ON
02/22/2009 12:30:00 CHANNEL 02 HIRS Electronics ON
02/22/2009 12:30:00 CHANNEL 05 HIRS Electronics ON
02/22/2009 12:30:00 CHANNEL 12 HIRS Electronics ON
02/22/2009 12:30:00 CHANNEL 13 HIRS Electronics ON
02/22/2009 12:30:00 CHANNEL 14 HIRS Electronics ON
02/22/2009 12:30:00 CHANNEL 16 HIRS Electronics ON
02/22/2009 12:30:00 CHANNEL 18 HIRS Electronics ON
02/22/2009 12:30:00 CHANNEL 19 HIRS Electronics ON
02/22/2009 12:30:00 CHANNEL 20 HIRS Electronics ON
02/22/2009 12:30:00 ELECTRONICS HIRS Electronics ON
02/22/2009 12:30:00 FILTER WHEEL HIRS Electronics ON
02/22/2009 12:30:00 SCAN MOTOR HIRS Electronics ON
02/22/2009 12:30:00 CHANNEL 01 HIRS Electronics ON
02/22/2009 12:30:00 CHANNEL 15 HIRS Electronics ON